# Variational Inference in Location-Scale Families: Exact Recovery of the Mean and Correlation Matrix

### v2

@@ 2.3 Symmetries and Bayesian inference
Figure 1: Posterior distribution of $\beta_{1}$ for a Bayesian logistic regression with $N$ examples. Vertical lines indicate the means estimated by MCMC and VI. These estimates match when the posterior is symmetric $(N\!=\!0,128)$ and differ when it is not $(N\!=\!4)$.
@@
[…]
@@
where $\boldsymbol{\beta}=(\beta_{0},\beta_{1},\beta_{2})$ and ${\bf x}_{i}=(1,x_{i1},x_{i2})$ is the $i^{\text{th}}$ component of a design matrix with two covariates. We compare the posteriors estimated by ADVI and, as a benchmark, long runs of Markov chain Monte Carlo (MCMC) for different numbers of examples. Both algorithms are implemented in the statistical software Stan [Carpenter et al., 2017]; see Appendix C.1 for details of these algorithms and Appendix C.2 for additional experimental results. Figure 1 plots these estimated posterior distributions for $\beta_{1}$ and $N\!=\!0,4,128$ examples. We see that the true posterior, as reported by MCMC, is even-symmetric for $N\!=\!0$ (where the posterior reduces to the prior) and very nearly symmetric for $N\!=\!128$, but clearly asymmetric for $N\!=\!4$, a regime where neither the prior nor the likelihood dominates.
Fig. 1 illustrates another result: the posterior mean is accurately estimated by VI when the posterior is symmetric ($N\!=\!0$ and $N\!=\!128$), but less accurately estimated when it is not ($N\!=\!4$). We now formalize this result, providing guarantees for VI when the target density is even-symmetric.
@@ 3.2 Illustrative examples
Figure 2: Robustness of VI to misspecifications. Left: VI with a factorized Gaussian, $q$, exactly recovers the mean of a multivariate student-t, $p$. Right: VI with a univariate Gaussian recovers the point of symmetry in target densities with different tails. This point equals the mean when the target is Laplace or student-t, but not when it is Cauchy (whose mean does not exist).
@@
Figure 2 (left) illustrates how VI behaves when $\mathcal{Q}$ is the family of Gaussians with diagonal covariance matrices and $p$ is a heavy-tailed two-dimensional student-t with $k\!=\!10$ degrees of freedom and correlation 0.9 between $z_{1}$ and $z_{2}$. Despite these misspecifications, $p$ and $\mathcal{Q}$ satisfy all the conditions of the theorem, and VI exactly recovers the mean of $p$.
@@
Figure 2 (right) illustrates how VI behaves when $\mathcal{Q}$ is the family of one-dimensional Gaussians and $p$ is (i) a Laplace distribution, (ii) a student-t distribution with $k\!=\!10$ degrees of freedom, and (iii) a Cauchy distribution. In all of these cases, VI recovers the point of symmetry of $p$; this point is equal to the mean of $p$ for the Laplace and student-t distributions, and to the median of $p$ for the Cauchy distribution (whose mean does not exist).
Figure 3: Approximating the mixture $p$ of two normals in eq. (15) when the modes are close ($m\!=\!1$) or well separated ($m\!=\!10$). Left: probability densities. Right: KL divergence between $p$ and a single normal $q_{\nu}$ with mean $\nu$. Per Theorem 8, $\text{KL}(q_{\nu}||p)$ has a stationary point at $\nu\!=\!0$. This stationary point is a minimizer when $m\!=\!1$ but a local maximizer when $m\!=\!10$.
@@
$p(z)=0.5\ \mathcal{N}(z;-m,1)+0.5\ \mathcal{N}(z;m,1),$ (15)
In this example, $p$ is symmetric about the origin, but it is not log-concave, thus violating one of the theorem’s conditions. The figure illustrates two cases—one where the modes of $p$ are close, with $m\!=\!1$ in eq. (15), and another, where they are well separated, with $m\!=\!10$. Per the theorem, $\text{KL}(q_{\nu}||p)$ has a stationary point at the origin. However, when $p$ has well-separated modes, this stationary point is a not a local minimizer of $\text{KL}(q||p)$; instead it is a local maximizer. In this case, VI would not center its approximation at the origin, but instead at one of the two modes.
@@ 4.1 Theoretical statement
$\vspace{3ex}-\,\int\!\log p_{0}\big{(}M^{-\frac{1}{2}}S^{\frac{1}{2}}\zeta\big{)}\,q_{0}(\zeta)\,\text{d}\zeta.$ (19)
@@
$\int\text{d}\zeta_{\backslash\{i\}}\int\text{d}\zeta_{i}\,\Big{[}f^{\prime}(\gamma\|\zeta\|)\,g(\|\zeta\|)\,\|\zeta\|^{-1}\Big{]}\zeta_{j}\zeta_{i}$ (26)
and note that the bracketed term in the integrand is spherically symmetric (and hence also invariant under the change of variables $\zeta_{i}\rightarrow-\zeta_{i}$). Now if $i\!\neq\!j$, then the integrand as a whole is an odd-symmetric function of $\zeta_{i}$, and hence the integral over all positive and negative values of $\zeta_{i}$ vanishes. On the other hand, if $i\!=\!j$, then from the spherical symmetry of the bracketed term it follows that the integral has the same value for all $i$. This implies, in turn, that all the diagonal elements on the right side of eq. (25) are equal.
@@
where we leveraged the fact the surface area of a $d$-dimensional hypersphere of radius $r$ is $2\pi^{\frac{d}{2}}r^{d-1}/\Gamma\left(d/2\right)$. Note that the left side of eq. (28) decreases monotonically with $\gamma$ from infinity to zero; thus the right side will determine a unique solution if it is a positive non-decreasing function of $\gamma$. Now, the log concavity of $p$ implies that the derivative of $\log p$ is monotone. Then, it follows from eqs. (18) and (22) that $-f^{\prime}(\gamma r)$ is a non-decreasing function of $\gamma$. It must also be true that $-f^{\prime}(0)\geq 0$, since otherwise $\log p$ would have a non-concave kink at the origin. Thus the right side of eq. (28) is indeed a positive non-decreasing function of $\gamma$, and a unique solution for $\gamma$ is determined. ∎
@@ 4.2 Illustrative examples
where $k$ denotes the number of degrees of freedom in the Student-t target, $M$ its scale matrix, and $S$ the covariance matrix of the Gaussian approximation. In this example, the elements of $M$ were one along the diagonal and constant off the diagonal ($S_{ij}=0.9$ for $i\neq j$). The results in the figure were obtained by ADVI for target densities with heavier tails ($202\!\leq\!k\!\leq\!$ 20) than $q$, particularly for small values of $k$. The results verify Theorem 10: $S$ matches $M$ up to a multiplicative constant and thus yields the exact correlation matrix of $p$.
@@ 5 NUMERICAL EXPERIMENTS
Table 1: Scaled errors ($\Delta_{\text{mean}},\Delta_{\text{corr}},\Delta_{\text{cov}}$) from VI when estimating the mean, correlation, and covariance of the target density, $p$. Errors are averaged across all elements of mean vectors and correlation/covariance matrices, and data sets are ordered by the degree to which $p$ violates the assumption of even symmetry, as measured by $\varepsilon_{90}$ in eq. (32). As a trend, less symmetric targets yield worse estimates of the mean and correlation; also, the correlation can be well-estimated even when the covariance is not.
For each distribution in Table  1, we use a stochastic procedure to assess whether it is approximately even-symmetric. Given a draw $z$ from $p(z)$—obtained by MCMC for the more complex distributions—we compute the reflected sample $z^{\prime}=2\mu\!-\!z$, where $\mu$ is the (estimated) mean of $z$ with respect to the density $p$. For each $(z,z^{\prime})$ pair, we then compute the violation
@@
If $\pi(z|x)$ is even-symmetric about $\mu$, then $\varepsilon(z)$ vanishes for all $z$. We evaluate $\varepsilon(z)$ at many (MCMC) samples and report ${}^{\text{th}}$ quantile $\varepsilon_{90}$ in Table 1.
@@ Acknowledgments
We thank four anonymous referees for their constructive feedback.
@@ Appendix A SYMMETRY OF LIKELIHOOD
A natural condition for the likelihood to be symmetric is for the data itself to be symmetric. To illustrate this, we generate $N$ observations. First we create $N=3$ adversarial samples $x=(-1,0,20)$. These observations are deliberately asymmetric, and produce asymmetric likelihoods (Figure 8). Next we generate $N=3$ ideal samples, $x=(-1,0,1)$, which are symmetric and produce symmetric likelihoods. Finally, we randomly draw $N=10$ and $N=100$ observations from a symmetric distribution $p$ (somewhat arbitrarily, we chose $p$ to be Gaussian though any well-behaved symmetric distribution would do). While no particular realization of the data is exactly symmetric, the data is “on average” symmetric. Already for $N=10$, we find that all non-Gaussian likelihoods are approximately symmetric, with the symmetry becoming stronger for $N=1000$.
@@ Appendix B SUPPORTING PROOFS
We can define another open set $\Omega_{\lambda}$ in $\mathbb{R}^{d}$ by translating the points of $\Omega$ by an amount $-\nu_{\lambda}$; namely,
@@ C.2 Additional results for Bayesian logistic regression
Figure 10 plots the marginal posterior densities for the Bayesian logistic regression introduced in section 2.3 with $N\!=\!0,1,2,4,8,16,32,128$ observations. (For clarity, the main body only shows the marginal posteriors for $\beta_{1}$ with $N\!=\!0,4,128$. From the figure, we see how the asymmetry manifests for all regression parameters as the number of examples is increased from $N\!=\!0$ (where the prior dominates) to $N\!=\!128$) (where the likelihood dominates). When the posterior is asymmetric, the posterior mean can still be accurately estimated (e.g., $N\!=\!32$), and in many cases we see that one component (but not all three) of the posterior mean is recovered.
@@ C.3 Targets in numerical experiments
  • $N$, the number of surveyed Peregrine falcons,
@@
[…]
@@
$\nu_{\mathrm{local}}$ $=$ 11 11
$\nu_{\mathrm{global}}$ $=$ 11 11
$s_{\mathrm{slab}}$ $=$ 22 22
$s_{\mathrm{df}}$ $=$ 100100 100100
$c_{0}$ $=$ $5$ (65)

### v3

@@ 2.3 Symmetries and Bayesian inference
Figure 1: Posterior distribution of $\beta_{1}$ for a Bayesian logistic regression with NN examples. Vertical lines indicate the means estimated by MCMC and VI. These estimates match when the posterior is symmetric (N=0,128)(N\!=\!0,128) and differ when it is not (N=4)(N\!=\!4).
@@
$\beta_{0},\beta_{1},\beta_{2}$ $\overset{\text{iid}}{\sim}$ $\text{Laplace}(0,0.5),$
@@
where $\boldsymbol{\beta}=(\beta_{0},\beta_{1},\beta_{2})$ and ${\bf x}_{i}=(1,x_{i1},x_{i2})$ is the $i^{\text{th}}$ component of a design matrix with two covariates. We compare the posteriors estimated by ADVI and, as a benchmark, long runs of Markov chain Monte Carlo (MCMC) for different numbers of examples. Both algorithms are implemented in the statistical software Stan [Carpenter et al., 2017]; see Appendix C.1 for details of these algorithms and Appendix C.2 for additional experimental results. Figure 1 plots these estimated posterior distributions for $\beta_{1}$ and N=0,4,128N\!=\!0,4,128 examples. We see that the true posterior, as reported by MCMC, is even-symmetric for N=0N\!=\!0 (where the posterior reduces to the prior) and very nearly symmetric for N=128N\!=\!128, but clearly asymmetric for N=4N\!=\!4, a regime where neither the prior nor the likelihood dominates.
Fig. 1 illustrates another result: the posterior mean is accurately estimated by VI when the posterior is symmetric (N=0N\!=\!0 and N=128N\!=\!128), but less accurately estimated when it is not (N=4N\!=\!4). We now formalize this result, providing guarantees for VI when the target density is even-symmetric.
@@ 3.2 Illustrative examples
Figure 2: Robustness of VI to misspecifications. Left: VI with a factorized Gaussian, $q$, exactly recovers the mean of a multivariate student-t, pp. Right: VI with a univariate Gaussian recovers the point of symmetry in target densities with different tails. This point equals the mean when the target is Laplace or student-t, but not when it is Cauchy (whose mean does not exist).
@@
Figure 2 (left) illustrates how VI behaves when $\mathcal{Q}$ is the family of Gaussians with diagonal covariance matrices and $p$ is a heavy-tailed two-dimensional student-t with k=10k\!=\!10 degrees of freedom and correlation 0.9 between $z_{1}$ and $z_{2}$. Despite these misspecifications, $p$ and $\mathcal{Q}$ satisfy all the conditions of the theorem, and VI exactly recovers the mean of $p$.
@@
Figure 2 (right) illustrates how VI behaves when $\mathcal{Q}$ is the family of one-dimensional Gaussians and $p$ is (i) a Laplace distribution, (ii) a student-t distribution with k=10k\!=\!10 degrees of freedom, and (iii) a Cauchy distribution. In all of these cases, VI recovers the point of symmetry of $p$; this point is equal to the mean of $p$ for the Laplace and student-t distributions, and to the median of $p$ for the Cauchy distribution (whose mean does not exist).
Figure 3: Approximating the mixture $p$ of two normals in eq. (15) when the modes are close (m=1m\!=\!1) or well separated (m=10m\!=\!10). Left: probability densities. Right: KL divergence between $p$ and a single normal $q_{\nu}$ with mean $\nu$. Per Theorem 8, $\text{KL}(q_{\nu}||p)$ has a stationary point at $\nu\!=\!0$. This stationary point is a minimizer when m=1m\!=\!1 but a local maximizer when m=10m\!=\!10.
@@
[…]
In this example, $p$ is symmetric about the origin, but it is not log-concave, thus violating one of the theorem’s conditions. The figure illustrates two cases—one where the modes of $p$ are close, with m=1m\!=\!1 in eq. (15), and another, where they are well separated, with m=10m\!=\!10. Per the theorem, $\text{KL}(q_{\nu}||p)$ has a stationary point at the origin. However, when $p$ has well-separated modes, this stationary point is a not a local minimizer of $\text{KL}(q||p)$; instead it is a local maximizer. In this case, VI would not center its approximation at the origin, but instead at one of the two modes.
@@ 4.1 Theoretical statement
$\vskip 12.91663pt-\,\int\!\log p_{0}\big(M^{-\frac{1}{2}}S^{\frac{1}{2}}\zeta\big)\,q_{0}(\zeta)\,\text{d}\zeta.$ (19)
@@
[…]
and note that the bracketed term in the integrand is spherically symmetric (and hence also invariant under the change of variables $\zeta_{i}\rightarrow-\zeta_{i}$). Now if $i\!\neq\!j$, then the integrand as a whole is an odd-symmetric function of $\zeta_{i}$, and hence the integral over all positive and negative values of $\zeta_{i}$ vanishes. On the other hand, if i=ji\!=\!j, then from the spherical symmetry of the bracketed term it follows that the integral has the same value for all $i$. This implies, in turn, that all the diagonal elements on the right side of eq. (25) are equal.
@@
where we leveraged the fact the surface area of a $d$-dimensional hypersphere of radius $r$ is $2\pi^{\frac{d}{2}}r^{d-1}/\Gamma\left(d/2\right)$. Note that the left side of eq. (28) decreases monotonically with $\gamma$ from infinity to zero; thus the right side will determine a unique solution if it is a positive non-decreasing function of $\gamma$. Now, the log concavity of $p$ implies that the derivative of $\log p$ is monotone. Then, it follows from eqs. (18) and (22) that $f^{\prime}(\gamma r)$ is a non-decreasing function of $\gamma$. It must also be true that $f^{\prime}(0)\geq 0$, since otherwise $\log p$ would have a non-concave kink at the origin. Thus the right side of eq. (28) is indeed a positive non-decreasing function of $\gamma$, and a unique solution for $\gamma$ is determined. ∎
@@ 4.2 Illustrative examples
where $k$ denotes the number of degrees of freedom in the Student-t target, $M$ its scale matrix, and $S$ the covariance matrix of the Gaussian approximation. In this example, the elements of $M$ were one along the diagonal and constant off the diagonal ($S_{ij}=0.9$ for $i\neq j$). The results in the figure were obtained by ADVI for target densities with heavier tails ($2\!\leq\!k\!\leq\!20$) than $q$, particularly for small values of $k$. The results verify Theorem 10: $S$ matches $M$ up to a multiplicative constant and thus yields the exact correlation matrix of $p$.
@@ 5 NUMERICAL EXPERIMENTS
Table 1: Scaled errors ($\Delta_{\text{mean}},\Delta_{\text{corr}},\Delta_{\text{cov}}$) from VI when estimating the mean, correlation, and covariance of the target density, pp. Errors are averaged across all elements of mean vectors and correlation/covariance matrices, and data sets are ordered by the degree to which $p$ violates the assumption of even symmetry, as measured by $\varepsilon_{90}$ in eq. (32). As a trend, less symmetric targets yield worse estimates of the mean and correlation; also, the correlation can be well-estimated even when the covariance is not.
For each distribution in Table  1, we use a stochastic procedure to assess whether it is approximately even-symmetric. Given a draw $z$ from p(z)p(z)—obtained by MCMC for the more complex distributions—we compute the reflected sample $z^{\prime}=2\mu\!-\!z$, where $\mu$ is the (estimated) mean of $z$ with respect to the density $p$. For each $(z,z^{\prime})$ pair, we then compute the violation
@@
[…]
@@ Acknowledgments
We thank four anonymous referees for their constructive feedback and Isaac Rankin for additional feedback on our manuscript.
@@ Appendix A SYMMETRY OF LIKELIHOOD
A natural condition for the likelihood to be symmetric is for the data itself to be symmetric. To illustrate this, we generate $N$ observations. First we create $N=3$ adversarial samples x=(1,0,20)x=(-1,0,20). These observations are deliberately asymmetric, and produce asymmetric likelihoods (Figure 8). Next we generate $N=3$ ideal samples, x=(1,0,1)x=(-1,0,1), which are symmetric and produce symmetric likelihoods. Finally, we randomly draw $N=10$ and $N=100$ observations from a symmetric distribution $p$ (somewhat arbitrarily, we chose $p$ to be Gaussian though any well-behaved symmetric distribution would do). While no particular realization of the data is exactly symmetric, the data is “on average” symmetric. Already for $N=10$, we find that all non-Gaussian likelihoods are approximately symmetric, with the symmetry becoming stronger for $N=1000$.
@@ Appendix B SUPPORTING PROOFS
We can define another open set $\Omega_{\lambda}$ in $\mathbb{R}^{d}$ by translating the points of $\Omega$ by an amount $\nu_{\lambda}$; namely,
@@ C.2 Additional results for Bayesian logistic regression
Figure 10 plots the marginal posterior densities for the Bayesian logistic regression introduced in section 2.3 with N=0,1,2,4,8,16,32,128N\!=\!0,1,2,4,8,16,32,128 observations. (For clarity, the main body only shows the marginal posteriors for $\beta_{1}$ with N=0,4,128N\!=\!0,4,128. From the figure, we see how the asymmetry manifests for all regression parameters as the number of examples is increased from N=0N\!=\!0 (where the prior dominates) to N=128N\!=\!128) (where the likelihood dominates). When the posterior is asymmetric, the posterior mean can still be accurately estimated (e.g., N=32N\!=\!32), and in many cases we see that one component (but not all three) of the posterior mean is recovered.
@@ C.3 Targets in numerical experiments
  • NN, the number of surveyed Peregrine falcons,
@@
$z_{1},z_{2}\overset{\text{iid}}{\sim}\frac{1}{2}\mathcal{N}(-1,2)+\frac{1}{2}\mathcal{N}(3,1).$ (64)
@@
$\nu_{\mathrm{local}}$ $=$ $1$
$\nu_{\mathrm{global}}$ $=$ $1$
$s_{\mathrm{slab}}$ $=$ $2$
$s_{\mathrm{df}}$ $=$ $100$
$c_{0}$ $=$ $5.$ (65)